OpenStudy (anonymous):

ok.... -2(7r+3)+6(7r+5)

OpenStudy (cwrw238):

simplify?

OpenStudy (anonymous):

yes simplify the expression by combining like terms

OpenStudy (cwrw238):

first distribute -2 over the first bracket and then +6 over the second bracket then combine like terms

OpenStudy (anonymous):

ah.. ok

OpenStudy (cwrw238):

i'll start you off -2(7r + 3) = -14r - 6

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

so the second part id 42r-30?

OpenStudy (anonymous):

is

OpenStudy (cwrw238):

42r + 30

OpenStudy (cwrw238):

so now you need to simplify -14r - 6 + 42r + 30

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for the first part, you divide by -42?

OpenStudy (anonymous):

err 14r sorry

OpenStudy (cwrw238):

no there is only addition and subtraction involved here a simple example is -x - 2 + 2x + 4 first bring like terms together:- -x + 2x - 2 + 4 = x + 2

OpenStudy (cwrw238):

so in your example the firat step is -14r + 42r - 6 + 30 now you work out the 2 terms with r in them and then the 2 numbers